# Manipulation of Quantum Strings by Edge Defect in Kagome Rydberg Atom Array

### v1

@@ Abstract
Optical tweezers can grab particles with their laser beam fingers, so these tiny point objects can be moved by using the laser light. However, manipulation of topological defects, especially higher-dimension ones becomes extremely difficult because they usually require non-local operation involving a great deal of operators, e.g. string operator. Here, using a large-scale quantum Monte Carlo method, we demonstrate the quantum string excitation can be grabbed in the Kagome Rydberg atom array with well-designed edges. After inserting an edge defect, the fractional charge at one end of the string is found to be pinned, allowing the quantum string to be manipulated. Since the quantum string is directed, its quantum fluctuations can be clearly observed due to the movement of the dangling fractional charge at the other end. Then, we can drag the two ends of the string in opposite directions, which is analogous to the separation of a quark-anti-quark pair connected by a gluon string. As the separation distance increases, the quantum fluctuations of the string are strongly suppressed. Eventually, a string-breaking phenomenon occurs. However, if both edge defects are located on the same side, the strings are always broken because both dangling fractional charges take the same gauge charge and can not annihilate each other into the vacuum. This work not only proposes an ideal experimental platform for analyzing quantum string excitations in the ground state but also provides a much simpler method for manipulating quantum string excitations in frustrated systems, such as spin ice.
@@
Introduction— Field lines can depict the fundamental interactions between particles, e.g., electric field lines between a positron and an electron and the gluon string between quarks. On the other hand, due to the interplay between geometry frustration and the strong interaction between spins, a line-type structure, also known as a ’string,’ can emerge in quantum frustrated magnetismFulde and Pollmann (2008); Jiang and Emig (2005); Zhou et al. (2022a, 2023); Zhang et al. (2016); Liu et al. (2024); Xiong (2023); Morris et al. (2009); Castelnovo et al. (2008); Mengotti et al. (2011); Wan and Tchernyshyov (2012). Interactions between the strings can result in various exotic topological phases of matter, such as the incommensurate phase Zhang et al. (2016); Zhou et al. (2023), “Luttinger string liquid” Zhou et al. (2022a), and U(1) lattice gauge theory (LGT) Zhou et al. (2022b); Zhang et al. (2018). Unlike the classical Dirac string in spin ice Castelnovo et al. (2008); Morris et al. (2009), here the strings are quantum, meaning the ground state of the string is a superposition of all possible fluctuating string structures, which can be considered as the worldsheet in 2+1 dimensions. Although the quantum string can be indirectly observed by analyzing the dynamical spectrum Zhou et al. (2022a); Liu et al. (2024), direct detection is still challenging.
@@
As a typical frustrated system, the Kagome lattice is composed of a corner-shared structure Han et al. (2012); Damle and Senthil (2006); Isakov et al. (2006); He et al. (2014); He and Chen (2015); He et al. (2015); Zhang and Eggert (2013); Zhang et al. (2018); Liu et al. (2024). The antiferromagnetic Ising interaction can impose a geometric constraint analogous to the ice rule on the system Pauling (1935), so that the disordered ground state preserves macroscopic degeneracy. If quantum fluctuations are introduced, higher-order perturbative interactions can lift the degeneracy, so the system enters into the valence bond solid phase Moessner et al. (2000); Isakov and Moessner (2003); Damle and Senthil (2006); Isakov et al. (2006), which breaks the translational symmetry. Due to the strong local constraint, the elementary excitations should be fractional charges with quantum strings attached Zhang and Eggert (2013); Liu et al. (2024). Both fractional charges and quantum strings are found to play a critical role in deconfined quantum criticality Senthil et al. (2004a, b); Liu et al. (2024), and its low-energy physics can be understood by LGT He and Chen (2015); He et al. (2015); Zhang et al. (2018); Liu et al. (2024), which involves the U(1) lattice gauge field strongly coupled to matter fields. Thanks to the rapid development of the Rydberg atom array Browaeys and Lahaye (2020); Nogrette et al. (2014); Bernien et al. (2017); Semeghini et al. (2021); Ebadi et al. (2021); Scholl et al. (2021), the Kagome geometry can be easily implemented. The ground state of the Kagome Rydberg atom array is still under debate Samajdar et al. (2021); Yan et al. (2023a), but the flexibility of the geometry and in situ manipulation make it possible to prepare classical string structures in the experiment Gonzalez-Cuadra et al. (2024). However, searching for the manipulation method of quantum strings in the ground state with more flexibility is still a critical task.
@@
In this manuscript, we study the Kagome Rydberg atom array with open boundary conditions mimicking the real experiment. Due to the long-range repulsive forces, the ground state becomes the stripe phase in which the fractional charges are confined. As shown in Fig. 1, after kicking out one site at the boundary, a “directed” quantum string spontaneously emerges with one side attached to the edge defect. Then, by further kicking out another site on the opposite side, both ends of the quantum string are tweezed, so that we can easily stretch the string until it breaks. The fractional charges at both broken ends take different charges, just like the positive-negative electron pair creation from the vacuum. Before breaking, the quantum string exhibits the asymptotic freedom-like behavior.
@@
Model— The system we consider is the Kagome Rydberg atom array, in which two-level atoms are individually trapped by optical tweezer lights. In each site, the ground state can be coupled to the Rydberg state via a two-photon process with site-independent tunable Rabi frequency and detuning . Atoms in the Rydberg state feel the long-range repulsive interaction with a Van der Waals (VdW) type form , and the strength of can be adjusted by changing the lattice spacing. Then, the Hamiltonian can be expressed as follows:
@@
where denotes the exciting process between the ground state and Rydberg state and is the density operator of atom in Rydberg state. The model above can be mapped into the extended quantum transverse Ising model by implementing the following transformation . Then, the VdW potential term plays the role of Ising interaction, so the Rydberg blockade effect imposes the local constraint. Meanwhile, the Rabi frequency and detuning correspond to the transverse and longitudinal magnetic field. When taking with lattice space set to one, the leading contribution of the diagonal term can be rewritten as which indicates only one atom in each triangle can be excited to the Rydberg state. Such a strong local constraint is called the triangle rule which causes the macroscopic degeneracy. However, if further long-range repulsive interactions are considered, the degeneracy can also be lifted so that the ground state becomes the stripe phase in which the atoms in one of three sublattices are fully excited (see Fig.1).
@@
Quantum String—To interpret the excitation of fractional charges and quantum strings, it is better to first introduce the lattice gauge mapping in brief. Same as the spin-1/2 XXZ model Zhang and Eggert (2013); He et al. (2014); He and Chen (2015); He et al. (2015); Zhang et al. (2018); Liu et al. (2024), as shown in Fig. 1(a), the “electric field” can be defined as at the bisector of the corner-shared triangle, where labels the triangles that compose the honeycomb dual-lattice. The positive direction is set as pointing from the down triangle to the up triangle. Then, we can immediately find that the triangle rule is indeed the “Gauss’s law”, and the violated configurations have different charges calculated by summing over all field lines around one triangle (as listed in Fig. 1(a)). Therefore, there are four different fractional excitations (gauge charge) classified by positive or negative charge and up- or down-triangle. Meanwhile, their gap energies are approximately equal at and can be tuned by changing the detuning strength Liu et al. (2024). Furthermore, after mapping, we can find that the stripe phase is very special because it represents a uniform “electric field” in the language of LGT. To explicitly observe the quantum string, we should take it as the reference vacuum state. As shown in Fig. 1(c), after subtracted by the reference state, the non-zero electric field forms a single string with zero electric field in the background.
@@
In the spin-1/2 XXZ model, the spin exchange or XY interaction provides the quantum fluctuation. In the LGT, it represents the production of the gauge charge pair or the hopping of the gauge charge at the next nearest neighbor dual-sites. In comparison, as demonstrated in Fig. 1(b), the Rabi term has the same effects but at the nearest neighbor site in the dual lattice. Both interactions keep the charge conservation. On the other hand, a sixth-order perturbation can contribute to the ring exchange interaction () in each hexagon without breaking the triangle rule Moessner et al. (2000); Isakov and Moessner (2003). Therefore, the string can gain kinetic energy from the movement of the gauge charge at the end and also from the fluctuation of the main body.
@@
As presented in Fig. 1(c), if we kick out one site at one side, it is equivalent to inserting an empty down-triangle or negative gauge charge . Then, it can move into the bulk with a string attached, but the other end of the string is always pinned at the edge defect, just like the dog leash. As the string length increases, the stripe order is disrupted more severely. Therefore, the string has a tension energy due to the long-range interactions. If the Rabi frequency is small, the fluctuation energy of the string cannot overcome the tension energy, so the gauge charge is confined. In contrast, with a larger Rabi frequency, deconfinement is expected. In the following numerical simulation, we set to ensure that the gauge charges are confined in the bulk, and the string attaching to the edge defect can emerge with a finite length at the same time.
@@
Results—The QMC method we utilized is the newly developed high-performance large-scale loop algorithm Xu and Zhang (2024), which can help achieve a large system size. To mimic the real experiment, we take open boundary conditions in both directions. The shape of the edge is designed as shown in Fig. 1(c) so that the outermost sites are fully occupied due to the edge effect. The nearest-neighbor repulsive interaction and the detuning are set to one, and the VdW long-range interactions are truncated to the third nearest-neighbor site. The maximum system size achieved reaches atom array sites. The temperature is set to , so the ground state physics can be observed. All the results are calculated by taking the average of 128 independent Markov chains with samples in each.
@@
Without the edge defect, the ground state is a perfect stripe phase that can be detected by measuring the local density of the atom in the Rydberg state . In the stripe phase, all the atoms in the C sublattice are excited to the Rydberg state, accompanied by local quantum fluctuations. Then, we can turn off one edge site by changing the local detuning to a large negative value to ensure the atom always remains in state . To highlight the quantum string excitation, we subtract it by the reference stripe state, which means the density in the C sublattice is transformed as . Furthermore, to illustrate the dynamics of the string, we also calculate the density of the resonant hexagon Zhang and Eggert (2013); Zhang et al. (2018), which characterizes the strength of the ring exchange process.
@@
As demonstrated in Fig. 2(a), the densities of both Rydberg excitation and resonant hexagon are higher around the edge defect. The interesting feature is that the sites in the blue dashed lines labeled region have apparent high intensity, although the strength decays as the site is further away from the edge defect. To figure out the physical reason, we make a snapshot, which is similar to taking a picture of the Rydberg atom distribution in the experiment. From Fig. 2(b), we can immediately observe the string with finite length attached to the edge. Meanwhile, we find that the string is directed, which means it cannot turn around to form a loop-like structure. The same phenomena are also observed in the XXZ model Zhang and Eggert (2013). The main reason is that the string can only move up left or up right. In addition, we can find some spots in other places, and their appearance is due to local quantum fluctuations from the vacuum which reflects the gauge charge confinement.
@@
At the other end of the quantum string, a gauge charge randomly moves in the middle of the bulk. Then, we can imagine that if an additional gauge charge appears, they will annihilate each other if they have opposite charges. Therefore, we turn off one site on the other side to produce another string with the gauge charge having the opposite charge. Remarkably, as shown in Fig. 3(a), we can only observe a single string with two ends attached to different edge defects. This strongly supports our expectation that the introduced gauge charge annihilates the existing gauge charge in the bulk. Consequently, the two strings merge into a single string. Then, both ends of the string are pinned, and the string vibration can be clearly observed in the bulk, which resembles the vibrating strings in a guitar or the ancient Chinese instrument–“Guzheng”.
@@
If we change the position of both edge defects, the string is stretched. As demonstrated in Fig. 3(b-c), the fluctuation region continuously shrinks along with the increase of the distance between defects. Nevertheless, its shape always maintains a rectangular form. Such characteristics can be understood from Fig. 3(d-f). When both ends of the string are fixed, its length is also fixed because the string is directed, or no two segments of the string lie in the same row. Then, the configuration with the fewest resonant hexagons can be used to define the boundary of the fluctuation region, e.g., the left boundary shown in Fig. 3(d-f). Consequently, it is easy to deduce how the string accommodates a longer defect distance.
@@
When the defect distance increases to the width of the system , the string is straightened, as shown in Fig. 4(a). The resonant fluctuation densities appear higher along the string, but this may result from local quantum fluctuations because the string cannot deform, as demonstrated in Fig. 4(c). Furthermore, once the defect distance is further enlarged, the string is overstretched. Then, we can observe that it breaks into two strings attached to different sides, as explicitly shown in Fig. 4(b).
@@
The string-breaking can be considered a sudden change in the system, so it should be reflected in the energies. Thus, we calculate the energies with different defect distances . Here, to maintain the left-right symmetry of the system, the defect distance can only take even numbers. Then, we obtain the energy difference and plot it in Fig. 4(d). Before the string-breaking occurs, the energy cost is continuously increasing while stretching the string, which implies that the quantum string is more free when edge defects are close to each other. Such phenomena are reminiscent of the counterintuitive property of the strong force, which exhibits asymptotic freedom in quark confinement.
@@
As a one-dimensional quantum object, the quantum string can usually be well described by a one-dimensional quantum model, such as in triangular lattices Zhang et al. (2016); Zhou et al. (2023), Kagome lattices Zhang and Eggert (2013); Liu et al. (2024), and even the high-Tc problem Wang et al. (2024). Here, the configuration of the string can be mapped into an effective spin-1/2 chain by mapping the gauge field () to effective spin up (down) Zhang and Eggert (2013). Then, the ring exchange interaction plays the role of the effective spin-exchange interaction. Although the second and third-order long-range repulsive interactions contribute to the complicated Ising interaction, the ring exchange is dominant. Consequently, the effective model describing the string is approximately the spin-1/2 XY model, which can be exactly solved by applying the Jordan-Wigner transformation. Counting the spins in Fig. 3(d-f), we can find that the defect distance can change the total magnetization as . Therefore, zero magnetization corresponds to and also to half-filled Jordan-Wigner fermions, which preserve the lowest energy. Then, stretching the string is equivalent to inserting additional Jordan-Wigner free fermions into a higher Fermi energy band, so the energy cost becomes larger and larger. Thus, it exhibits asymptotic freedom-like behavior.
Furthermore, we also attempt to place both edge defects on the same side in the numerical simulation. Then, we observed each defect is attached to a string. If the defects are close to each other, the fluctuation of the strings becomes asymmetric, which implies there are interactions between the strings. These two strings cannot merge into one string because the fractional charges at the free ends of both strings take the same charge.
Conclusion and Outlook—We successfully pin the quantum string at the boundary by introducing an edge defect in the Kagome Rydberg atom array, allowing us to manipulate it effectively. With the help of an additional edge defect, both ends of the string can be grabbed tightly, enabling us to stretch this one-dimensional quantum curve smoothly. Although the entire process is not dynamic but simulated in the ground state, it can still be considered as following a quasi-static path. After measuring the density distribution of the Rydberg excitation and resonant hexagons, we find that the string is directed, and its fluctuation is confined to a rectangular region. Finally, by examining the variation of the energy, we observe asymptotic freedom phenomena before string-breaking occurs, which can be well understood by the free Jordan-Wigner fermionic theory.
As mentioned before, in the real experiment, it is easy to kick out a site by changing the detuning. Definitely, the tweezer light at that site can also be switched off. However, considering the possibility of dynamically moving the string step by step in future experiments, the previous scheme would be better. In addition, annealing to the ground state would not be easy due to the topological features of the string, so searching for a better annealing method would also be an interesting direction Yan et al. (2023b); Ding et al. (2024).
@@
Note on related work—During the preparation of this manuscript, we noticed that a related experimental work observed the string-breaking phenomena in the Kagome Rydberg atom array using a different experimental scheme Gonzalez-Cuadra et al. (2024). In comparison, our work focuses on ground-state physics with ten times larger system size, and study the interplay between gauge charge and quantum string by the manipulation of it.
@@
Acknowledgement– X.-F. Z. acknowledges funding from the National Science Foundation of China under Grants No. 12274046, No. 11874094, No.12147102 and No.12347101, Chongqing Natural Science Foundation under Grants No. CSTB2022NSCQ-JQX0018, Fundamental Research Funds for the Central Universities Grant No. 2021CDJZYJH-003, and Xiaomi Foundation / Xiaomi Young Talents Program.

### v2

# Geometric Breaking of Quantum Strings in Kagome Rydberg Atom Array

@@ Abstract
Leveraging the rapid development of quantum simulators, the intriguing phenomena of string breaking are observed across various quantum simulation platforms Gonzalez-Cuadra et al. (2024); Cochran et al. (2024). However, the strings can only be broken by quenching the tension energy due to the system size limitation, such ”tensile breaking” behaves quite the same as in one-dimension De et al. (2024). To discover the distinctive properties of quantum strings in two dimensions, we utilize a newly developed quantum Monte Carlo method to simulate a ten times larger scale (743-site) Kagome Rydberg atom array which can be described with U(1) lattice gauge theory. With designed edges, one end of the string can be easily grasped, allowing for clearly observing the dynamics of the dangling gauge charge at the open end. Then, we can drag each end in opposite directions perpendicular to the string. As the separation distance increases, asymptotic freedom-like behavior is detected. Eventually, the string breaks because the distance between both ends surpasses the string’s length, so named ”geometric breaking”. Distinct from ”tensile breaking,” the two newly formed strings are suddenly uncorrelated, accompanied by a sudden energy increase. Lastly, we study multiple strings and successfully prepare a heart-like superposition state of broken and unbroken quantum strings with three gauge charges attached reminiscent of quarks and gluon. This work can benefit the comprehension of quantum strings and also shed light on the simulation of high-energy physics.
@@
Introduction— Field lines can depict the fundamental interactions between particles, e.g., electric field lines between a positron and an electron and the gluon string between quarks. On the other hand, due to the interplay between geometry frustration and the strong interaction between spins, a line-type structure, also known as a ’string,’ can emerge in quantum frustrated magnetism Fulde and Pollmann (2008); Jiang and Emig (2005); Zhou et al. (2022a, 2023); Zhang et al. (2016); Liu et al. (2024); Xiong (2023); Morris et al. (2009); Castelnovo et al. (2008); Mengotti et al. (2011); Wan and Tchernyshyov (2012). Interactions between the strings can result in various exotic topological phases of matter, such as the incommensurate phase Zhang et al. (2016); Zhou et al. (2023), “Luttinger string liquid” Zhou et al. (2022a), and U(1) lattice gauge theory (LGT) Zhou et al. (2022b); Zhang et al. (2018); Liu et al. (2024). Unlike the classical Dirac string in spin ice Castelnovo et al. (2008); Morris et al. (2009), here the strings are quantum, meaning the ground state is a superposition of all possible fluctuating string structures, which can be considered as a worldsheet in 2+1 dimensions.
@@
As a typical frustrated system, the Kagome lattice is composed of a corner-shared structure Han et al. (2012); Damle and Senthil (2006); Isakov et al. (2006); He et al. (2014); He and Chen (2015); He et al. (2015); Zhang and Eggert (2013); Zhang et al. (2018); Liu et al. (2024). The antiferromagnetic Ising interaction can impose a geometric constraint analogous to the ice rule on the system Pauling (1935), so that the disordered ground state preserves macroscopic degeneracy. If quantum fluctuations are introduced, higher-order perturbative interactions can lift the degeneracy, so the system enters into the valence bond solid phase Moessner et al. (2000); Isakov and Moessner (2003); Damle and Senthil (2006); Isakov et al. (2006). Due to the strong local constraint, the elementary excitations should be fractional charges with quantum strings attached Zhang and Eggert (2013); Liu et al. (2024). Both fractional charges and quantum strings are found to play a critical role in deconfined quantum criticality Senthil et al. (2004a, b); Liu et al. (2024), and its low-energy physics can be understood by LGT He and Chen (2015); He et al. (2015); Zhang et al. (2018); Liu et al. (2024), which involves the U(1) lattice gauge field strongly coupled to matter fields. Thanks to the rapid development of the Rydberg atom array Browaeys and Lahaye (2020); Nogrette et al. (2014); Bernien et al. (2017); Semeghini et al. (2021); Ebadi et al. (2021); Scholl et al. (2021), the Kagome geometry can be easily implemented Gonzalez-Cuadra et al. (2024). The ground state of the Kagome Rydberg atom array is still under debate Samajdar et al. (2021); Yan et al. (2023), but the flexibility of the geometry and in-situ manipulation makes it possible to prepare the string excitation Gonzalez-Cuadra et al. (2024). However, the small system size (59 atoms) in the experiment constrains that the string can only be broken by fixing both ends and adjusting the tension energy. As a straightforward way, deforming the quantum string by moving the ends in both and directions becomes unfeasible in recent experiment conditions. Furthermore, as mentioned in the outlook of Ref.Gonzalez-Cuadra et al. (2024), analyzing the interplay between multiple quantum strings becomes an even more complex task which is vital in higher dimensions.
@@
In this manuscript, we numerically simulate a large-size Kagome Rydberg atom array with open boundary conditions mimicking the real experiment. As shown in Fig. 1, after kicking out one site at the boundary, a “directed” quantum string spontaneously emerges with one side attached to the edge defect. Then, by further introducing another edge defect on the opposite side, both ends of the quantum string are tweezed, so that we can easily stretch the string until it breaks. Before this geometric breaking happens, the quantum string exhibits the asymptotic freedom-like behavior. At last, we discuss the interaction between multiple strings and the superposition of broken and unbroken quantum strings caused by vacuum fluctuation.
@@
Model— The Kagome Rydberg atom array is composed of the two-level atoms individually trapped by optical tweezer lights. The ground state can be coupled to the Rydberg state via a two-photon process with site-independent tunable Rabi frequency and detuning . Atoms in the Rydberg state feel the long-range repulsive interaction with a Van der Waals (VdW) type form , and the strength of can be adjusted by changing the lattice spacing. Then, the Hamiltonian can be expressed as follows:
@@
where denotes the exciting process and is the density operator of atom in Rydberg state. The model above can be mapped into the extended quantum transverse Ising model by implementing the following transformation . Then, the VdW potential term plays the role of Ising interaction, and the Rabi frequency and detuning correspond to the transverse and longitudinal magnetic field. When taking with lattice space set to one, the leading contribution of the diagonal term can be rewritten as which indicates only one atom in each triangle can be excited to the Rydberg state. If further long-range repulsive interactions are considered, the degeneracy can be lifted so that the ground state becomes the stripe phase in which the atoms in one of three sublattices are fully excited (see Fig.1).
@@
Quantum String—It is better to first briefly introduce the lattice gauge theory before discussing quantum string. Same as the spin-1/2 XXZ model Zhang and Eggert (2013); He et al. (2014); He and Chen (2015); He et al. (2015); Zhang et al. (2018); Liu et al. (2024), as shown in Fig. 1(a), the “electric field” is defined as at the bisector of the corner-shared triangle, where labels the triangles that compose the honeycomb dual-lattice. The positive direction is set as pointing from the down triangle to the up triangle. Then, we can immediately find that the triangle rule is indeed the “Gauss’s law”, and the violated configurations have different charges calculated by summing over all field lines around one triangle (as listed in Fig. 1(a)). Therefore, there are four different fractional excitations (gauge charges or matter fields) classified by positive or negative charge and up- or down-triangle. Meanwhile, their gap energies are approximately equal at and can be tuned by changing the detuning strength Liu et al. (2024). To explicitly observe the quantum string, we should take the stripe phase as the reference vacuum state. As shown in Fig. 1(c), after subtracted by the reference state, the non-zero electric field forms a single string with zero electric field in the background.
@@
In the spin-1/2 XXZ model, the spin exchange or XY interaction provides the quantum fluctuation. In the LGT, it represents the production of the gauge charge pair or the hopping of the gauge charge at the next nearest neighbor dual-sites. In comparison, as demonstrated in Fig. 1(b), the Rabi term has the same effects but at the nearest neighbor site in the dual lattice. Both interactions keep the charge conservation. On the other hand, a sixth-order perturbation can contribute to the ring exchange interaction () in each hexagon without breaking the triangle rule Moessner et al. (2000); Isakov and Moessner (2003). Therefore, the string can gain kinetic energy from the movement of the gauge charge and also from the quantum fluctuation of the gauge field.
@@
As presented in Fig. 1(c), removing one lower-edge site is equivalent to inserting a negative gauge charge (e.g. ). Then, it can hop into the bulk with a string attached, but the other end of the string is always pinned at the edge defect, just like the dog leash. As the string length increases, the stripe order is disrupted more severely accompanied by the energy cost strongly increases, which means the string feels a tension energy caused by the long-range interactions. When the Rabi frequency is small, the fluctuation energy of the string cannot overcome the tension energy, so the gauge charge is confined. In contrast, deconfinement is expected at a larger Rabi frequency. In the following, we set to make sure the gauge charges are confined in the bulk, and the string attaching to the edge defect can emerge with a finite length at the same time.
@@
Results—The QMC method we utilized is the newly developed high-performance large-scale algorithm Xu and Zhang (2024), and the open boundary conditions in both directions are taken to mimic the real experiment. The shape of the edge is designed as shown in Fig. 1(c) so that the outermost sites are fully occupied due to the edge effect. The nearest-neighbor repulsive interaction and the detuning are set to one, and the VdW long-range interactions are truncated to the third nearest-neighbor site which most of the simulation taken Yan et al. (2023); Gonzalez-Cuadra et al. (2024). We define the length of the first row in sublattice C as , the number of rows as , so Fig. 1(c) has sites with and . The maximum system size achieved reaches atom array sites, ten times larger than the state-of-the-art experiment Gonzalez-Cuadra et al. (2024). The temperature is set to to approach the ground state physics. All the results are calculated by taking the average of 128 independent Markov chains with samples in each.
@@
In the stripe phase, all the atoms in the C sublattice are excited to the Rydberg state, accompanied by local quantum fluctuations. Then, we can turn off one edge site by changing the local detuning to a large negative value to ensure the atom always stays in state . To highlight the quantum string excitation, we subtract it by the reference stripe state, which means the density in the C sublattice is transformed as . Furthermore, to illustrate the dynamics of the string, we also calculate the density of the resonant hexagon Zhang and Eggert (2013); Zhang et al. (2018), which characterizes the strength of the ring exchange process. In all figures, the intensity of the resonant configuration is normalized by the maximum value.
@@
As demonstrated in Fig. 2(a), the densities of both Rydberg excitation and resonant hexagon are higher around the edge defect. The interesting feature is that the sites inside the blue dashed lines labeled region have apparent high intensity, and the strength decays as the site is further away from the edge defect. To figure out the physical reason, we make a snapshot of the transformed Rydberg state distribution. From Fig. 2(b), we can immediately observe the string with a finite length attached to the edge. Meanwhile, we find that the string is directed along the -direction, which means it cannot turn around to form a loop-like structure. Such mobility constraint of quasi-particle is also observed in the XXZ model Zhang and Eggert (2013) and widely exists in the fracton systems Gromov and Radzihovsky (2024); Zhou et al. (2021); Wiedmann et al. (2024). In addition, we can find some spots in other places, and their appearance is due to local quantum fluctuations from the vacuum reminiscent of the Schwinger mechanism.
@@
At the open end of the quantum string, a gauge charge randomly moves in the middle of the bulk. Then, we can imagine that if an additional gauge charge appears, they will annihilate each other if they take opposite charges. Therefore, we turn off one site on the other side to produce another string with the gauge charge taking the opposite charge. Remarkably, as shown in Fig. 3(a), we can only observe a single string with two ends attached to different edge defects. This strongly supports our expectation that the introduced gauge charge annihilates the existing gauge charge in the bulk. Consequently, the two strings merge into a single string. Then, both ends of the string are pinned, and the string’s transverse vibration can be observed, which resembles the vibrating classical strings in a violin or the ancient Chinese instrument–“Guzheng”.
@@
If we change the position of both edge defects, the string is stretched in -direction. The distance between defects in -direction can be defined as . As demonstrated in Fig. 3(b-c), the fluctuation region continuously shrinks along with the increase of defect distance . Nevertheless, its shape always maintains a rectangular form. Such characteristics can be understood from Fig. 3(e). When both ends of the string are static, the string length is fixed to (numbers of arrows and ) because the string is directed and no two segments of the string lie in the same row. Then, the classical configuration with the fewest possible resonant hexagons can be used to define the boundary of the fluctuation region, e.g., the string in Fig. 3(e) marks the most left boundary. Consequently, it is easy to deduce how the string accommodates a longer defect distance.
@@
When the defect distance increases to half of the string length, the string is straightened, as shown in Fig. 3(c). The resonant fluctuation densities appear higher along the string, but this may result from local quantum fluctuations because the string cannot deform, as demonstrated in Fig. 3(e). Then, once the defect distance is further enlarged, the string is overstretched due to geometric reasons. Then, we can observe that it breaks into two strings attached to different sides, as explicitly shown in Fig. 3(d).
@@
Different from the “tensile breaking”, this “geometric breaking” is a sudden change in the system which can be reflected in the energies. Thus, we calculate the energies with different defect distances . Here, to maintain the left-right symmetry of the system, the defect distance can only take even numbers. Then, we obtain the energy difference and plot it in Fig. 3(f). Before the string-breaking occurs, the energy cost is continuously increasing while stretching the string, which implies that the quantum string is more free when edge defects are close to each other. Such phenomena are reminiscent of the counterintuitive property of the strong force, which exhibits asymptotic freedom in quark confinement.
@@
As a one-dimensional quantum object, the quantum string can be well described by a one-dimensional quantum model, such as in triangular lattices Zhang et al. (2016); Zhou et al. (2023), Kagome lattices Zhang and Eggert (2013); Liu et al. (2024), and even the high-Tc problem Wang et al. (2024). Here, the configuration of the string can be mapped into an effective spin-1/2 chain by mapping the gauge field () to effective spin up (down) Zhang and Eggert (2013). Then, the ring exchange interaction plays the role of the effective spin-exchange interaction. Although the second and third-order long-range repulsive interactions contribute to the complicated Ising interaction, the ring exchange is dominant while stretching. Consequently, the effective model describing the string is approximately the spin-1/2 XY model, which can be exactly solved by applying the Jordan-Wigner transformation. Counting the spins in Fig. 3(e), we can find that the defect distance can change the total magnetization of effective spins as . Therefore, zero magnetization corresponds to and also to half-filled Jordan-Wigner fermions, which preserve the lowest energy. Then, stretching the string is equivalent to inserting additional Jordan-Wigner free fermions into a higher Fermi energy band, so the energy cost becomes larger and larger which exhibits asymptotic freedom-like behavior.
Furthermore, we also consider the stretching in both - and -direction. As shown in Fig.4(a), although the defect distance is not large enough to trigger the “geometric breaking”, the features of breaking can still be observed from density distribution when stretching string in direction by enlarging . This phenomenon is related to the “tensile breaking” in which the tension energy (O()) is larger than the gap energy (O()) of gauge charges. However, different from stretching in the direction, we can still observe a blurry string between defects while breaking happens. It means the ground state may be a superposition of broken and unbroken status. Therefore, the energy differences do not show any sudden jump in Fig.4(b) which is strongly different from the quench dynamics observed in Ref.Cochran et al. (2024); Gonzalez-Cuadra et al. (2024).
At last, we prefer to discuss the interplay between multiple strings. The reverse process of geometric breaking can be considered as two strings with opposite gauge charge move closely and merge into single one. Differently, if the static gauge charges of both strings are set on the same side, they can not merge because the gauge charge on the open ends takes the same charge. Meanwhile, they will feel an effective repulsive interaction between them due to no-crossing conditions.
Interestingly as depicted in Fig.4(c), when setting three defects on different sides, the single quantum string attached to one side will be frustrated to choose one of two quantum strings on the other side to merge. Then, the ground state can be constructed as a superposition of broken and unbroken quantum strings with the help of the vacuum fluctuation in the middle bulk. Consequently, as demonstrated in Fig.4(d), the heart-like structure emerges. Meanwhile, the strong intensities of the resonant configuration in the middle region due to intensive breaking-merging processes support the strong vacuum fluctuations.
Conclusion and Outlook—We successfully manipulate the quantum string by introducing an edge defect in the Kagome Rydberg atom array. Then both ends of the string can be grabbed tightly in large system size, enabling us to adiabatically stretch it in a perpendicular direction. Strongly different from most discussed “tensile breaking” in one and two dimensions, an asymptotic freedom-like behavior is observed before sudden “geometric breaking” happens when the string is overstretched. At last, we demonstrate a superposition of broken and unbroken strings attached with three static gauge charges that can be prepared by introducing three defects on the edges. In the real experiment, the edge site can be removed by adjusting the detuning or directly switching off the tweezer light at that site. However, considering the possibility of dynamically moving the string step by step in future experiments, the previous scheme would be recommended.
@@
Note on related work—During the preparation of this manuscript, we noticed that a related experimental work observed the string-breaking phenomena in the Kagome Rydberg atom array using a different experimental scheme Gonzalez-Cuadra et al. (2024). In comparison, our work focuses on ground-state physics with more than ten times larger system size, so that the strings can be stretched in diversified ways. Meanwhile, the interplay between multiple strings is also considered. Furthermore, our strategy of manipulating quantum string is more simple because only one site needs to be engineered. The main conclusion of our work can be immediately verified in their experimental platform with a large system size.
@@
Acknowledgement– X.-F. Z. acknowledges funding from the National Science Foundation of China under Grants No. 12274046, No. 11874094, No.12147102, and No.12347101, Chongqing Natural Science Foundation under Grants No. CSTB2022NSCQ-JQX0018, Fundamental Research Funds for the Central Universities Grant No. 2021CDJZYJH-003, and Xiaomi Foundation / Xiaomi Young Talents Program.